OpenStudy (xenesis98):

You have a bag with 4 pairs of socks. There are 2 red pairs and 2 black pairs. You pull a sock from the bag and put it back 3 times. What is the probability that you pull out a red sock each time?

OpenStudy (therawbugeyes):

hey so i neeed the answers please

OpenStudy (xenesis98):

a.1/6 b.1/8 c.1/4 d.1/2 e.-1/10

OpenStudy (peachpi):

2 out of 4 pairs are red, so what's the probability that you pick a red pair the first time?

OpenStudy (xenesis98):

1/2?

OpenStudy (xenesis98):

or 50%

OpenStudy (peachpi):

right. Then you put it back, for each pick so the probability that you pick a red sock is 1/2 for the second and third picks as well.

OpenStudy (xenesis98):

confused

OpenStudy (xenesis98):

so is 1/2 the answer?

OpenStudy (anonymous):

50% chance every time

OpenStudy (xenesis98):

so its right? lol

OpenStudy (anonymous):

Yes

OpenStudy (peachpi):

sorry. The page glitched on me. It's a 1/2 chance all three times, so you multiply them together to get the overall probability. \[\frac{ 1}{ 2 }\times \frac{ 1 }{ 2 }\times \frac{ 1 }{ 2 }=\frac{ 1 }{ 8 }\]